OpenStudy (anonymous):

What is the simplified form of 3 over 4x plus 3 + 21 over 8 x squared minus 14x minus 15 ? @SolomonZelman @midhun.madhu1987 @FriedRice14

OpenStudy (anonymous):

@anonymous_user

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

\[ \frac{3}{4x+3}+\frac{21}{8x^2-14x-15}\]? ?

OpenStudy (anonymous):

Right

OpenStudy (kirbykirby):

WHat you could do is factor the denominator of the 2nd fraction: \(8x^2 -14x-15 = 8x^2+6x-20x-15 = 2x(4x+3)-5(4x+3)=(4x+3)(2x-5)\)

OpenStudy (kirbykirby):

The last thing is just (2x-5), just the parenthesis just got cutt off

OpenStudy (anonymous):

that's it?

OpenStudy (kirbykirby):

No that's one step

OpenStudy (kirbykirby):

You can then combine the 2 fractions into one by multiplying the top and bottom of the first fraction by (2x-5): \[ \frac{3}{4x+3}+\frac{21}{(4x+3)(2x-5)}\\ ~ \\=\frac{3}{(4x+3)}\frac{(2x-5)}{(2x-5)}+\frac{21}{(4x+3)(2x-5)}\\~\\=\frac{3(2x-5)+21}{(4x+3)(2x-5)} \\ ~ \\ =\frac{6x-15+21}{(4x+3)(2x-5)} \\ ~ \\ \frac{6(x+1)}{(4x+3)(2x-5)}\]

OpenStudy (kirbykirby):

now i'm not sure if your teacher wants you to keep the top and bottom as factors or expanded out..

OpenStudy (anonymous):

This is good thanks

OpenStudy (kirbykirby):

yw!